Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
bersanetti, boldrini, casanueva, lunghini, mantovani, menzione, nardecchia, pinto, spinicelli, tomelleri - 17:07 Friday 31 October 2025 (68076) Print this report
DAS tuning to reduce the MICH offset

This morning, starting from the standard TCS configuration used throughout the run, we began exploring a possible DAS tuning aimed at reducing the MICH offset, which is currently around 70 (compared to ~20 measured before). A total of three steps were performed, all involving an increase of the inner ring power on the NI.

STEP 1
NI inner ring power increased by 20%.
NI PICKOFF value: 0.578 W → 0.583 W.
Action completed at 08:44 UTC.

STEP 2
NI inner ring power further increased by 20%.
NI PICKOFF value: 0.583 W → 0.589 W.
Action completed at 09:55 UTC.

At this point, we manually unlocked the ITF to test the lock acquisition.
The ITF relocked at the first attempt at CARM NULL 1F at 11:13 UTC, and, after imposing a MICH offset of about 70, it relocked at LN3 at 12:02 UTC.

STEP 3
NI inner ring power further increased by 20%.
NI PICKOFF value: 0.589 W → 0.603 W.
Action completed at 12:54 UTC.

After this step, some glitches were observed in several signals, including the BNS range (see the attached figure), suggesting a possible degradation of ITF stability. For this reason, we decided to revert the third step.
Consequently, at 14:36 UTC, I reduced the NI inner ring power back to the value of STEP 2.

Then, the ITF unlocked spontaneously at 14:50 UTC and relocked at CARM NULL 1F on the first attempt at 15:09 UTC.
The ITF reached LN3 at 15.57 UTC. 

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×